# Natal, Brazil

### v1

@@ 3 Methodology
Participation was voluntary and the estimated time to complete the survey was 10 to 15 minutes. A total of 500 emails were sent out over a period of 45 days (January 30 to March 13, 2023), and 28 responses were received, resulting in a response rate of 5.6%.
We generate a customized survey for each of the 121 final projects with the evolution of the project’s specific CI practices. To accomplish this, we calculate the values of seven CI practices that would be presented to the developers, as follows:
  • Commit Per Weekday CITheater2019 : Mean of the absolute number of commits according to the weekday of the analyzed period.
  • Coverage CITheater2019 : It measures which parts of a program are executed when running the tests. Represents the percentage of the program covered by tests.
  • Build Duration CITheater2019 : It measures the duration of the build (build finished at timestamp - build started at timestamp).
  • Build Activity/Build-frequency STAHL201448 : It is a unit interval (i.e., a closed interval [0,1]) representing the rate of builds across days, i.e., if builds were made every day in the period of CI, the value would be 1. If builds were made in half of the days, the value would be 0.5. If there were no builds, the value would be 0.
  • Build Health/Build Quality EliezioRevisaoSistematica2022 : It is a unit interval representing the rate of build failures across days. If there were build failures every day, the value would be 0. If there were no build failures, the value would be 1.
  • Time to Fix a Broken Build CITheater2019 : It consists of the median time in a period that builds remained broken. When a build breaks, we compute the time in seconds until the build returns to the “passed” status. If the CI period ends and the build does not return to the “passed” status, we consider the time since it was broken until the end of the period. When a period has no broken builds, the value would be 0.
  • Comments Per Change/Comments Per PR thompson2017large ; staahl2013experienced : Mean of the number of comments grouped by changes; normally, a change is a Pull Request when we are working at GitHub.
Before presenting the evolution charts for the seven CI practices we quantified for the projects, we outline the practices and describe how we measured them for the purpose of the survey.
@@
  • Base: The deployment process is manual or semi-manual with some parts scripted and rudimentarily documented in some way. The team will typically practice unit-testing and have one or more dedicated test environments
  • Beginner: Introduces frequent polling builds for faster feedback and build artifacts are archived for easier dependency management. The team starts to investigate ways of gradually automating the existing manual integration testing
@@
  • Advanced: Techniques for zero downtime deploys can be important to include in the automated process. Techniques of database migrations can be used to completely avoid manual routines for database updates. Fully automatic acceptance tests and maybe also generating structured acceptance criteria directly from requirements
@@
The survey consisted of 19 questions, 12 of which were open-ended. To give respondents as much flexibility as possible, 8 questions were mandatory (7 of which were close-ended while 1 was open-ended), and 11 questions were optional (all open-ended). The survey was divided into four sections:
@@
    Perception of CI practices: In this set of questions, we asked participants about the current state of CI in their projects. This included the level of CI maturity, what practices are under monitoring and how often, and which tools are used.
@@
    Evaluation of CI practice: In the last part of the survey, we presented evolution charts for the seven CI practices that we quantified for the project in question, as shown in Figure 2. We asked participants to rate the importance of these practices to the project and whether they deemed it important to measure the evolution of these practices over time. We also asked if the values presented for these practices were surprising or worrying in any way.
    For some projects, we presented correlations between CI practices and the number of closed pull requests (an indicator of effort) and the number of bug-related issues (an indicator of quality), as shown Figure 3. We requested participants to provide their comments regarding the correlations shown. Only significant correlations, above 50%, were shown. Otherwise, these correlations were omitted from participants.
    Finally, we asked if there was any missing information that our participants would like to know of or if they would like to receive regular updates. We also requested participants to provide their email should they wish to participate in the prize draw. We provide a complete copy of our questionnaire and dataset in our public repository. [^6] https://zenodo.org/records/10936371
Figure 2: Partial evolution charts for the seven CI practices
Figure 3: Partial correlations between the seven CI practices and quality attributes
We compile all survey results into a single spreadsheet. We show the results from the closed-ended responses in a graphical format, while we perform a Thematic Analysis of the open-ended responses, using quotes to illustrate the themes whenever necessary. We group the responses into seven themes, which are shown in Figure 4.
Figure 4: Survey Themes
@@
The steps involved in our study follow the guidelines from Linaker et al. linaker2015guidelines , which are: defining the research questions, defining the target audience, creating the survey, evaluating it on a subset of the sample population, conducting the survey, analyzing the results and drawing conclusions. We create a first draft of the survey and went through five rounds of revision with different participants drawn from our professional network (i.e., not the core developers yet), until we reached a refined version of the survey. We send this refined version to 20% ($100500\frac${100}{500}divide start_ARG 100 end_ARG start_ARG 500 end_ARG) of developers and awaited for the first responses (15 days). Once 6 responses were obtained, we made small adjustments to the questions, improving the clarity of our survey. We then send this final version to the remaining 80% ($400500\frac${400}{500}divide start_ARG 400 end_ARG start_ARG 500 end_ARG) of developers.
@@
Approach: In RQ3, we conduct a Gray Literature Review (GLR) garousi2020benefitting to compile the existing monitoring capabilities of existing CI tools and services. To systematically collate existing monitoring capabilities, we conduct a Document Analysis bowen2009documentanalysis on the official documentation of main CI services. In our analysis, we focused on the seven most popular CI services, covering 99% of all observed instances of CI usage GolzadehCIServer2022 : TravisCI, GHActions, CircleCI, AppVeyor, Azure, GitLab and Jenkins.
@@
To collect the official documentation from the CI services, we used the following query strings on google.com, using incognito windows:
  • (Travis) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
  • (“github actions”) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
  • (Jenkins) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
  • (GitLab) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
  • (CircleCI) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
  • (AppVeyor) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
  • (Azure) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
In this step, we ignore third-party documentation and sponsored links to ensure the trustworthiness of the data morgan2022conductingDocumentAnalysis . We examine additional 33 documents with this methodology. In Table 1 and Table 2, we list the documents found during the two steps of research in which the Thematic Analysis was applied. We provide Gray Literature URLs and store all data searched and collected in an external database [^7] https://zenodo.org/records/10936371 for later consultation to increase studies’ replicability KAMEI2021106609 .
@@
In the third step of RQ3, we decided to include the analysis of third-party tools, as these tools may have features not present in the main CI services. We collect all references to third-party tools cited in the official CI services documentation obtained from the previous two steps. Subsequently, we conduct a Document Analysis on the documentation of these third-party tools. In this step, we identify 31 mentions of third-party tools and conduct a complementary Document Analysis on 41 documents related to these tools. Similar to the first steps, we provided Gray Literature URLs and stored all the data searched and collected in an external database [^8] https://zenodo.org/records/10936371 for later consultation to enhance the replicability of the studies. In Table 3 and Table 4, we list all the documents analyzed.
@@ 31 SPLUNK manual snowballing
After sampling the texts, we store them in our database and conducted a Thematic Analysis maguire2017doingThematicAnalysis , where the aforementioned CI practices guided our process of generating themes. Therefore, the output of our Document Analysis is a set of themes that summarize the existing monitoring capabilities of CI tools in relation to CI practices (e.g., tools available to monitor code coverage levels). We identify 152 citations of CI practices related to monitoring practices across the 70 analyzed documents of CI services (as shown in Table 1 and Table 2 ) and 43 citations of CI practices related to monitoring practices across the 41 analyzed documents of third-party tools (as shown in Table 3 and Table 4 ). There is no standardization of CI practices among these various CI tools analyzed, since we have different nomenclatures that represent the same practice. To better organize our results, we conduct a standardization process for these CI practice citations to eliminate duplication, using the follow steps:
@@
    We eliminate variations in practice nomenclature across different CI services. For example, the citation “mean time to resolution” found in a GHActions document was standardized to “mean time to recovery”, which is the more prevalent term [^9] https://www.leanix.net/en/wiki/vsm/dora-metrics. Similarly, the citation “Lead time to change” was adjusted to “lead time for change”.
@@
    We also employ CI practices definitions, when available, to facilitate standardization. For instance, the Cicle CI documentation [^10] https://circleci.com/blog/how-to-measure-devops-success-4-key-metrics/ defines “Success Rate” as “the number of passing runs divided by the total number of runs over a period of time.” This definition closely aligns with the CI practice used in our survey: “Build Health”, hence we convert all citation of “Success Rate” to “Build Health”. Table 5 provides an overview of all standardization procedures conducted.
@@
    Subsequent to this process, we eliminate redundant CI practices citations, culminating in a final total of 32 unique themes.
Table 5: Standardization of CI Practice Citations
# CI practice citations standardized themes
@@ 33 Security checks $\rightarrow$ security vulnerabilities
Similarly to RQ1, during the analysis, we encountered practices beyond the scope of CI. Therefore, we classified these practices more generally as “DevOps practices”. Among them, the DORA practices were the most frequently mentioned in the official documentation of CI services. The definition of these practices are:
@@
Finally, as well as in RQ1, the extracted codes and themes were reviewed by another author. We also calculated the agreement score between the two sets of labels using the Cohen Kappa statistical coefficient cohen1968weighted , which measures the agreement between two “judges”. We obtained a score of 0.4502, indicating “Moderate Agreement”.
@@ 4 Results
The generated themes are shown in Figure 5. We group themes into 2 high-level themes: (i) “monitoring opportunities”, which contains mentions of CI practices but with no indication of monitoring those practices (e.g. there are recurrent build failures, but no indication as to whether developers know the current or tolerated level of build failures) and (ii) “monitored practices”, which contains evidence that some CI practices are monitored (e.g., automatic comments generated by the coverage analysis tools).
Figure 5: Themes that emerged from PRs comments
During the analysis, we find some practices outside the scope of CI, and we thus classify such practices under the more general “DevOps practices” umbrella.
We observe that developers are mainly monitoring 1) Coverage (e.g., “… will increase coverage by 0.02”), 2) Code quality (e.g., “SonarCloud Quality Gate failed… new-duplicated -lines-density”), and 3) security practices (e.g., “new alerts: 1 for Useless assignment to local variable”). Coverage was the most frequently monitored practice, with over 80% of themes ($258306\frac${258}{306}divide start_ARG 258 end_ARG start_ARG 306 end_ARG). CodeCov was the main tool used in most projects, responsible for the majority of coverage monitoring instances ($206258\frac${206}{258}divide start_ARG 206 end_ARG start_ARG 258 end_ARG).
@@
Our results suggest that there exist a concern with monitoring CI practices although these constitute a small minority $4324892\frac${432}{4892}divide start_ARG 432 end_ARG start_ARG 4892 end_ARG ( or 8%) of discussions held during the development of the analyzed projects. Whether this is an indication of lack of concern regarding monitoring cannot be answered by our methodological approach applied here in RQ1, since natural discussions through online comments cannot be used to measure the exact level of concern developers would have regarding a specific issue (e.g., monitoring a CI practice). As such, we contact developers in RQ2 to seek direct opinions regarding the importance of monitoring CI practices.
@@
The results of the survey questions are presented in this section, organized according to the themes generated through our Thematic Analysis.
@@ 9 CI reports 1
While participants mostly mention build and test criteria to define their CI maturity level, when asked which practices their projects monitor, respondents mentioned a variety of practices, as shown in Table 7. When asked to provide reference values to the CI practices in their projects (the most common values for these practices), only four practices were mentioned: code coverage, build duration, time to release a new version and commit frequency. Table 8 summarizes the mentioned values. In relation to frequency, Figure 8 shows that half of the participants $1428\frac${14}{28}divide start_ARG 14 end_ARG start_ARG 28 end_ARG state that monitoring is done daily or quarterly (every 3 months).
Table 7: Monitored practices mentioned by the developers
@@ 16 Vulnerability scan of the resulting docker image 1
Table 8: Reference values to the CI practices mentioned by developers
CI practice reference values
@@
Our responses suggest that some CI practices are not perceived as important to some participants. For example, Developer 1 expressed: “All of the above-mentioned items except ‘Commit code frequently.’ I prefer meaningful commits, not frequent ones.” Developer 25 stated: “All except 1 (why would commit frequency matter?).” Additionally, some participants do not see the value of monitoring in their projects. Developer 5 mentioned: “Most of these metrics are logged in GitHub, although we usually don’t analyze the trends or discuss the practices too much as a team, since many of the tracked values fluctuate (Avrae is a kind of monolith which leads to things like coverage being difficult to increase).”
We find evidence that projects with a lower level of CI (“Base” and “Beginner”), monitor less CI practices. When checking which CI practices these projects monitor, a maximum of 3 practices were mentioned. Developer 4: “We don’t actively monitor any of these.”, Developer 17: “Mainly 5.” and Developer 3: “build health.” Conversely, in projects deemed as more mature (“Expert” and “Advanced”), there was a tendency to monitor more practices: Developer 15: “code coverage, commit frequency, time of release of a new version, build duration, CVEs for security”, Developer 9: “1 (Commit code frequently), 3 (Fix broken builds immediately), 4 (Write automated developer tests), 5 (All tests and inspections must pass ), 7 (All tests and inspections must pass )”, Developer 12: “All of the above, except commit frequency and to some extent time of release of a new version (this rather depends on completion of the required new features for the version).”
Figure 9 shows the distribution of the number of practices mentioned by the developers. The data was organized into two distinct groups: (i) projects classified as “Expert and Advanced” and (ii) projects classified as “Base and Beginner”, excluding “Intermediate” projects. As observed, developers associated with the “Expert and Advanced” group, cited a higher number of monitored metrics.
Figure 9: Number of monitored practices cited by developers x Project Maturity
@@
Figure 10 shows an overview of the tools mentioned by participants to perform monitoring tasks in their projects. CodeCov is the most used by 42% ( or $126\frac{11}${26}divide start_ARG 11 end_ARG start_ARG 26 end_ARG ) of participants. In second place, GH-Actions itself is used by 30% (or $826\frac${8}{26}divide start_ARG 8 end_ARG start_ARG 26 end_ARG) of participants. GH-Actions automatically tracks basic build information like “build time” and “failed/passed builds,” which can be useful for monitoring. SonarCloud comes in third with 8% ( or $226\frac${2}{26}divide start_ARG 2 end_ARG start_ARG 26 end_ARG) of participants using the tool. Based on data from the currently used tools, Developer 20 summarizes the most common monitoring flow for CI practices: “Codecov for tracking code coverage automatically on each commit. GitHub Actions automatically tracks things like ’build time’, ’failed/passed builds’, etc.”, since CodeCov and GH-Actions are the most used tools.
@@
To sum up, most of participants monitor “code coverage” using CodeCov $121\frac{11}${21}divide start_ARG 11 end_ARG start_ARG 21 end_ARG (or 52% ) and information such as “build duration” and “build status” are monitored using GH-Actions $821\frac${8}{21}divide start_ARG 8 end_ARG start_ARG 21 end_ARG (or 38% ). Some developers monitor code quality practices using Sonarcloud $221\frac${2}{21}divide start_ARG 2 end_ARG start_ARG 21 end_ARG (or 9% ). Other tools are mentioned once (CDash, Trivy, Jenkins, sanitizers and Azure/container-scan). $421\frac${4}{21}divide start_ARG 4 end_ARG start_ARG 21 end_ARG (or 19% ) participants mentioned not using any tools to monitor CI practices. No tools were mentioned to monitor the remaining CI practices, such as “time to fix a broken build”, “build health” or “commit activity.” We noticed that there is a pressing concern regarding security and code quality in the CI pipeline. Developer 8 expressed: “Azure/container-scan GitHub Action: container best practices and vulnerabilities”, and Developer 12 said: “Codecov, SonarCloud, and sanitizers are part of the CI process and ensure that we don’t merge untested code to begin with, as part of the review process.”

### v2

@@ 3 Methodology
The survey consisted of 18 questions, 11 of which were open-ended. To give respondents as much flexibility as possible, 8 questions were mandatory (7 of which were close-ended while 1 was open-ended), and 10 questions were optional (all open-ended). Participation was voluntary and the estimated time to complete the survey was 10 to 15 minutes. A total of 500 emails were sent out over a period of 45 days (January 30 to March 13, 2023). We received 28 responses, all of which were considered valid as the participants completed the mandatory questions. Overall, our survey achieved a response rate of 5.6%. Across all 18 questions, participants answered an average of 15.25 questions, with a minimum of 8 (a single occurrence) and a maximum of 18 (nine occurrences). In relation to the number of responses received for each optional question, the optional question with the highest number of responses received 26 answers, while the least answered question received 15 responses with an average of 21 answers.
We generated a customized survey for each of the 121 final projects, detailing the evolution of the project’s specific CI practices. To achieve this, we calculated the values of seven CI metrics associated with these practices, which were presented to the developers who participated in the survey. Table1 presents the CI metrics used in the survey, along with their definitions and the related CI practices we want to evaluate.
Table 1: Continuous Integration Practices, Metrics and Definitions
CI Practice CI Metric CI Metric Definition
Commit Code Frequently Duvall2007 Commit Per Weekday CITheater2019
Mean of the absolute number of
commits according to the weekday
of the analyzed period.
Write Automated
Developer Tests
Duvall2007
Coverage CITheater2019
It measures which parts of a
program are executed
when running the tests.
Keep the Build Fast fowler2006continuous Build Duration CITheater2019
It measures the duration
of the build. The time between
the start and end of the build.
Automate the Build fowler2006continuous Build Activity STAHL201448
It is a unit interval
(i.e., a closed interval [0,1])
representing the rate of builds
across days.
Don’t Commit
Broken Code
Duvall2007 ; fowler2006continuous
Build Health EliezioRevisaoSistematica2022
It is a unit interval representing
the rate of build failures
across days.
Fix Broken Builds
Immediately
Duvall2007 ; fowler2006continuous
Time to Fix
a Broken Build CITheater2019
It consists of the median time
in a period that builds remained
broken.
Conduct Code Reviews mcintosh2016empirical Comments Per Change thompson2017large ; staahl2013experienced
Mean of the number of comments
grouped by changes.
Before presenting the evolution charts for the seven CI metrics we quantified for the projects, we outline the metrics and describe how we measured them for the purpose of the survey.
@@
  • Base: The deployment process is manual or semi-manual with some parts scripted and rudimentarily documented in some way. The team will typically practice unit-testing and have one or more dedicated test environments.
  • Beginner: Introduces frequent polling builds for faster feedback and build artifacts are archived for easier dependency management. The team starts to investigate ways of gradually automating the existing manual integration testing.
@@
  • Advanced: Techniques for zero downtime deploys can be important to include in the automated process. Techniques of database migrations can be used to completely avoid manual routines for database updates. Fully automatic acceptance tests and maybe also generating structured acceptance criteria directly from requirements.
@@
The survey was divided into four sections:
@@
    Perception of CI practices: In this set of questions, we asked participants about the current state of CI in their projects. This included the level of CI maturity, what metrics are under monitoring and how often, and which tools are used.
@@
    Evaluation of CI practice: In the last part of the survey, we presented evolution charts for the seven CI metrics that we quantified for the project in question, as shown in Figure 3. We asked participants to rate the importance of these metrics to the project and whether they deemed it important to measure the evolution of these metrics over time. We also asked if the values presented for these metrics were surprising or worrying in any way.
    For some projects, we presented correlations between CI metrics and the number of closed pull requests (an indicator of effort) and the number of bug-related issues (an indicator of quality), as shown Figure 4. We requested participants to provide their comments regarding the correlations shown. Only significant correlations, above 50%, were shown. Otherwise, these correlations were omitted from participants.
    Finally, we asked if there was any missing information that our participants would like to know of or if they would like to receive regular updates. We also requested participants to provide their email should they wish to participate in the prize draw. We provide a complete copy of our questionnaire and dataset in our public repository. [^8] https://doi.org/10.5281/zenodo.14569025
Figure 3: Partial evolution charts for the seven CI metrics
Figure 4: Partial correlations between the seven CI metrics and quality attributes
We compile all survey results into a single spreadsheet. We show the results from the closed-ended responses in a graphical format, while we perform a Thematic Analysis of the open-ended responses, using quotes to illustrate the themes whenever necessary. We group the responses into seven themes, which are listed below:
@@
The steps involved in our study follow the guidelines from Linaker et al. linaker2015guidelines , which were: defining the research questions, defining the target audience, creating the survey, evaluating it on a subset of the sample population, conducting the survey, analyzing the results and drawing conclusions. We created a first draft of the survey and went through five rounds of revision with different participants drawn from our professional network (i.e., not the core developers yet), until we reached a refined version of the survey. We sent this refined version to 20% ($100500\frac${100}{500}divide start_ARG 100 end_ARG start_ARG 500 end_ARG) of developers and awaited for the first responses (15 days). Once 6 responses were obtained, we made small adjustments to the questions, improving the clarity of our survey. We then sent this final version to the remaining 80% ($400500\frac${400}{500}divide start_ARG 400 end_ARG start_ARG 500 end_ARG) of developers.
@@
Approach: In RQ3, we conduct a Gray Literature Review (GLR) garousi2020benefitting to compile the existing monitoring capabilities of existing CI tools and services. Since this method enables faster acquisition of research results in practical contexts, it more accurately reflects the current capabilities of CI services and third-party tools in monitoring CI practices. To systematically collate existing monitoring capabilities, we conduct a Document Analysis bowen2009documentanalysis on the official documentation of main CI services. In our analysis, we focused on the seven most popular CI services, covering 99% of all observed instances of CI usage GolzadehCIServer2022 : TravisCI, GHActions, CircleCI, AppVeyor, Azure, GitLab and Jenkins.
@@
To collect the official documentation from the CI services, we used the following query string on google.com, using incognito windows:
  • (<CI Service Name>999¡CI Service Name¿ = “Travis”, “github actions”, “Jenkins”, “GitLab”, “CircleCI”, “AppVeyor”, “Azure”) AND (“continuous integration”) AND (practices OR metrics) AND monitoring
In this step, we ignore third-party documentation and sponsored links to ensure the trustworthiness of the data morgan2022conductingDocumentAnalysis . We examine additional 33 documents with this methodology. In Table 2 and Table 3, we list the documents found during the two steps of research in which the Thematic Analysis was applied. We provide Gray Literature URLs and store all searched and collected data in an external database [^10] https://doi.org/10.5281/zenodo.14569025 for later consultation to increase study’s replicability KAMEI2021106609 .
@@
In the third step of RQ3, we decided to include the analysis of third-party tools, as these tools may have features not present in the main CI services. We collect all references to third-party tools cited in the official CI services documentation obtained from the previous two steps. Subsequently, we conduct a Document Analysis on the documentation of these third-party tools. In this step, we identify 31 mentions of third-party tools and conduct a complementary Document Analysis on 41 documents related to these tools. Similar to the first steps, we provided Gray Literature URLs and stored all the data searched and collected in an external database [^11] https://doi.org/10.5281/zenodo.14569025 for later consultation to enhance the replicability of the study. In Table 4 and Table 5, we list all the documents analyzed.
@@ 31 SPLUNK manual snowballing
After sampling the texts, we store them in our database and conducted a Thematic Analysis maguire2017doingThematicAnalysis , where the aforementioned CI practices guided our process of generating themes. To identify CI practices, we examine the CI metrics monitored by these services.
The document analysis and thematic analysis were conducted to identify citations to CI metrics monitoring across the examined documents. The process involved open coding, performed manually by the first author, through a thorough examination of the citations of CI metrics. The core steps began with an initial open coding phase where the documents were read in detail, and codes were generated to represent the monitored metrics. We identify 152 citations (codes) of CI metrics related to monitoring practices across the 70 analyzed documents of CI services (as shown in Table 2 and Table 3 ) and 43 citations (codes) of CI metrics related to monitoring practices across the 41 analyzed documents of third-party tools (as shown in Table 4 and Table 5). After, a standardization process was applied to the identified CI metrics citations. This process involved the follow steps:
@@
    We eliminate variations in metric nomenclature across different CI services. For example, the citation “mean time to resolution” found in a GHActions document was standardized to “mean time to recovery”, which is the more prevalent term [^12] https://www.leanix.net/en/wiki/vsm/dora-metrics. Similarly, the citation “Lead time to change” was adjusted to “lead time for change”.
@@
    We also employ CI metrics definitions, when available, to facilitate standardization. For instance, the Cicle CI documentation [^13] https://circleci.com/blog/how-to-measure-devops-success-4-key-metrics/ defines “Success Rate” as “the number of passing runs divided by the total number of runs over a period of time.” This definition closely aligns with the CI practice used in our survey: “Build Health”, hence we convert all citation of “Success Rate” to “Build Health”. Table 6 provides an overview of all standardization procedures conducted.
@@
    Subsequent to this process, we eliminate redundant CI metrics citations.
These standardized metrics citations culminated in a final total of 32 unique themes, they were then grouped into high-level themes, with seven themes associated with CI services and nine themes linked to third-party tools, reflecting the tools where the metrics were observed.
Table 6: Standardization of CI metrics Citations
# CI metric citations standardized themes
@@ 33 Security checks $\rightarrow$ security vulnerabilities
Similarly to RQ1, during the analysis, we encountered metrics beyond the scope of CI. Therefore, we classified these practices more generally as “DevOps metrics”. Among them, the DORA metrics were the most frequently mentioned in the official documentation of CI services. The definition of these metrics are:
@@
Finally, as well as in RQ1, the extracted codes and themes were reviewed by another author. We also calculated the agreement score between the two sets of labels using the Cohen Kappa statistical coefficient cohen1968weighted , which measures the agreement between two “judges”. We obtained a score of 0.45, indicating “Moderate Agreement”. Due to the low value of the Cohen’s Kappa coefficient for the Document Analysis of RQ3, a third author reviewed the disagreements between the first author and the second author, with the final result being the code obtained from this second round of review. After the second round of review, we improved the Cohen Kappa statistical coefficient to 0.63, which indicates “Substantial agreement”.
@@ 4 Results
The themes generated for our Thematic Analysis, which highlight key patterns and insights derived from the data, are presented in Figure 5.
Figure 5: Themes that emerged from PR comments
During the analysis, we find some metrics outside the scope of CI, and we thus classify such practices under the more general “DevOps metrics” umbrella.
We observe that developers are mainly monitoring 1) Coverage (e.g., “… will increase coverage by 0.02”), 2) Code quality (e.g., “SonarCloud Quality Gate failed… new-duplicated -lines-density”), and 3) security practices (e.g., “new alerts: 1 for Useless assignment to local variable”). Coverage was the most frequently monitored metrics, with over 80% of themes ($258306\frac${258}{306}divide start_ARG 258 end_ARG start_ARG 306 end_ARG). CodeCov was the main tool used in most projects, responsible for the majority of coverage monitoring instances ($206258\frac${206}{258}divide start_ARG 206 end_ARG start_ARG 258 end_ARG).
@@
Our results suggest that there exists a concern with monitoring CI practices, although these constitute a small minority $4324892\frac${432}{4892}divide start_ARG 432 end_ARG start_ARG 4892 end_ARG ( or 8%) of discussions held during the development of the analyzed projects. Whether this is an indication of lack of concern regarding monitoring cannot be answered by our methodological approach applied here in RQ1, since natural discussions through online comments cannot be used to measure the exact level of concern developers would have regarding a specific issue (e.g., monitoring a CI practice). As such, we contact developers in RQ2 to seek direct opinions regarding the importance of monitoring CI practices.
@@
The results of the survey questions are presented in this section.
@@ 9 CI reports 1
While participants mostly mention build and test criteria to define their CI maturity level, when asked which practices their projects monitor, respondents mentioned a variety of practices/metrics, as shown in Table 8. When asked to provide reference values for the metrics corresponding to the monitored CI practices in their projects (the most common values for these metrics), only four metrics were mentioned: code coverage, build duration, time to release a new version and commit frequency. Table 9 summarizes the mentioned values. In relation to frequency, Figure 8 shows that half of the participants $1428\frac${14}{28}divide start_ARG 14 end_ARG start_ARG 28 end_ARG state that monitoring is done daily or quarterly (every 3 months).
Table 8: Monitored practices/metrics mentioned by the developers
@@ 16 Vulnerability scan of the resulting docker image 1
Table 9: Reference values to the CI metrics mentioned by developers
CI metric reference values
@@
Our responses suggest that some CI metrics are not perceived as important to some participants. For example, Developer 1 expressed: “All of the above-mentioned items except ‘Commit code frequently.’ I prefer meaningful commits, not frequent ones.” Developer 25 stated: “All except 1 (why would commit frequency matter?).” Additionally, some participants do not see the value of monitoring in their projects. Developer 5 mentioned: “Most of these metrics are logged in GitHub, although we usually don’t analyze the trends or discuss the practices too much as a team, since many of the tracked values fluctuate (Avrae is a kind of monolith which leads to things like coverage being difficult to increase).”
We find evidence that projects with a lower level of CI (“Base” and “Beginner”), monitor less CI metrics. When checking which CI metrics these projects monitor, a maximum of 3 metrics were mentioned. Developer 4: “We don’t actively monitor any of these.”, Developer 17: “Mainly 5.” and Developer 3: “build health.” Conversely, in projects deemed as more mature (“Expert” and “Advanced”), there was a tendency to monitor more metrics: Developer 15: “code coverage, commit frequency, time of release of a new version, build duration, CVEs for security”, Developer 9: “1 (Commit code frequently), 3 (Fix broken builds immediately), 4 (Write automated developer tests), 5 (All tests and inspections must pass ), 7 (All tests and inspections must pass )”, Developer 12: “All of the above, except commit frequency and to some extent time of release of a new version (this rather depends on completion of the required new features for the version).”
Figure 9 shows the distribution of the number of metrics mentioned by the developers. The data was organized into two distinct groups: (i) projects classified as “Expert and Advanced” and (ii) projects classified as “Base and Beginner”, excluding “Intermediate” projects. The intention of excluding intermediate projects was to specifically highlight the extreme groups where the effects of CI adoption can be more strongly perceived. As observed, developers associated with the “Expert and Advanced” group, cited a higher number of monitored metrics.
Figure 9: Number of monitored metrics cited by developers x Project Maturity
We applied statistical tests to compare the number of metrics between projects of two previous groups. Our sample for both groups are 16 entries. First, we conducted the Shapiro-Wilk test for normality, which indicated that the data for both groups did not significantly deviate from a normal distribution (p = 0.1494 for Expert-Advanced and p = 0.09315 for Base-Beginner). Levene’s test for homogeneity of variances further confirmed the variability of the data in each group is similar (p = 0.3575). Given these results, we performed an independent two-sample t-test, which revealed a statistically significant difference in the number of metrics used between the groups (p = 0.0027). The mean number of metrics for the Expert-Advanced group was 4.5, while the Base-Beginner group had a mean of 1.33, indicating that projects with higher CI maturity tend to use significantly more CI metrics.
@@
Figure 10 shows an overview of the tools mentioned by participants to perform monitoring tasks in their projects. CodeCov is the most used by 42% ( or $126\frac{11}${26}divide start_ARG 11 end_ARG start_ARG 26 end_ARG ) of participants. In second place, GHActions itself is used by 30% (or $826\frac${8}{26}divide start_ARG 8 end_ARG start_ARG 26 end_ARG) of participants. GHActions automatically tracks basic build information like “build time” and “failed/passed builds,” which can be useful for monitoring. SonarCloud comes in third with 8% ( or $226\frac${2}{26}divide start_ARG 2 end_ARG start_ARG 26 end_ARG) of participants using the tool. Based on data from the currently used tools, Developer 20 summarizes the most common monitoring flow for CI practices: “Codecov for tracking code coverage automatically on each commit. GitHub Actions automatically tracks things like ’build time’, ’failed/passed builds’, etc.”, since CodeCov and GHActions are the most used tools.
@@
To sum up, most of participants monitor “code coverage” using CodeCov $121\frac{11}${21}divide start_ARG 11 end_ARG start_ARG 21 end_ARG (or 52% ) and information such as “build duration” and “build status” are monitored using GHActions $821\frac${8}{21}divide start_ARG 8 end_ARG start_ARG 21 end_ARG (or 38% ). Some developers monitor code quality metrics using Sonarcloud $221\frac${2}{21}divide start_ARG 2 end_ARG start_ARG 21 end_ARG (or 9% ). Other tools are mentioned once (CDash, Trivy, Jenkins, sanitizers and Azure/container-scan). $421\frac${4}{21}divide start_ARG 4 end_ARG start_ARG 21 end_ARG (or 19% ) participants mentioned not using any tools to monitor CI metrics. No tools were mentioned to monitor the remaining CI metrics, such as “time to fix a broken build”, “build health” or “commit activity.” We noticed that there is a pressing concern regarding security and code quality in the CI pipeline. Developer 8 expressed: “Azure/container-scan GitHub Action: container best practices and vulnerabilities”, and Developer 12 said: “Codecov, SonarCloud, and sanitizers are part of the CI process and ensure that we don’t merge untested code to begin with, as part of the review process.”
GitHub provides several built-in quality and security tools, such as Dependabot and CodeQL, which are widely used for dependency health monitoring and vulnerability analysis. However, it is notable that these tools did not prominently appear in the survey responses. A plausible explanation is that these tools may have been perceived as part of the GitHub Actions ecosystem, which was mentioned by 30% of respondents. Since GitHub Actions integrates various CI functionalities, including dependency and security monitoring, respondents may have grouped these tools under the broader GitHub Actions umbrella. Another possibility is that the survey’s focus on “CI metrics” and it is possible that these tools were perceived as outside the scope of the CI process. This emphasis might have overshadowed the mention of tools like Dependabot and CodeQL in the response.